# A Trifecta of Modelling Tools: Navigating COMPAS, Shark and Bayesian Inference for Binary Black Hole Model Selection

### v1

@@ 2.2 Chemically homogeneous evolution
Mandel & de Mink (2016) found that the CHE channel significantly contributes to the BBH merger rate ($10.5\pm 0.5$ $\rm Gpc^{-3}yr^{-1}$ at redshift 0). These systems typically have a total mass ranging between $50-110$ $M_{\odot}$. Marchant et al. (2016) explores the parameter space for forming CHE binaries. They found that at low metallicity, BBHs will merge within the age of the Universe due to weaker winds ($t_{\rm delay,min}=0.4$ Gyr for $Z_{\odot}/50$). There is strong evidence for increasing chirp masses with decreasing metallicity, with masses above the pair instability SN gap. If these mergers were detected, they could probe the evolution of massive stars in the early Universe. In addition to calculating the merger rate and mass distribution for the CHE channel, du Buisson et al. (2020) finds that the cosmological merger rate ($5.8$ $\rm Gpc^{-3}yr^{-1}$ at redshift 0) can vary up to 40% for extreme deviations in the metallicity specific SFR.
In COMPAS, there are two CHE modes, ‘optimistic’ and ‘pessimistic’, which refer to how a CH star evolves on the MS (Riley et al., 2021). In the ‘optimistic’ mode, if the star begins the main sequence with fast enough rotation to evolve as a CHE star, it is assumed the star continues rotating fast enough to remain CH throughout the MS and the spin is not checked. For the CHE mode ‘pessimistic’ the spin is checked at every timestep, and if it has slowed below the CH metallicity-dependent rotational frequency threshold, it is immediately converted to a regular MS star and evolve to the end of the MS (Team COMPAS: Riley, J. et al., 2022).
@@ 2.3 Remnant mass prescription
GWs can be used to infer the underlying mass distributions of BBHs. This remnant mass distributions depend on the metallicity and the type of explosion mechanism. Their formation can be described in three stages; stellar collapse and bounce, convective engine, and post-explosion fallback. Massive stars undergo a series of successful burning stages, building an iron core until electron degeneracy pressure can no longer support this. Electron capture and dissociation of the core elements into alpha particles accelerates the core collapse. This collapse is halted when the core reaches nuclear densities and nuclear forces cause the core to bounce and emit a shock outwards. This results in a mixing of the hot and cold layers in the star, known as convection. The convective engine provides the means to convert potential energy released in the collapse into explosion energy in the form of a supernovae (Fryer et al., 2012; Fryer et al., 2021). The commonly accepted supernova explosion mechanism is neutrino driven, where some of the neutrinos emitted during the core collapse are reabsorbed, powering an explosion (Janka et al., 2007; Marek & Janka, 2009; Janka, 2012, 2017).
Fryer et al. (2012) provide analytic prescriptions for estimating the remnant mass in population synthesis codes, based on at the time recent understanding of supernova and gamma-ray burst explosions. In their model/fit, the remnant mass depends on the core, fallback mass and explosion mechanism. They explore two neutrino-driven explosion mechanisms– the rapid convection, which leads to explosion in the 250 ms after bounce, and the delayed convection, which can occur over a longer timescale. Fryer et al. (2022) altered the remnant mass prescription from Fryer et al. (2012) to study the growth rate of convection. They introduce a new parameter in their models denoted $f_{\rm mix}$, the mixing growth time for convective instabilities. A fast growth time, $f_{\rm mix}=4.0$, and a slow growth time, $f_{\rm mix}=0.5$, correspond to the rapid and delayed models respectively from Fryer et al. (2012). Whether a star undergoes a successful or failed supernovae provides clues on the final remnant object. However, there are uncertainties surrounding these mechanisms (Mapelli, 2021). GW detection from a core collapse SN can be distinguish between explosion mechanisms (Powell et al., 2016, 2023)
@@
Schneider et al. (2021) also applies the Mandel & Müller (2020) semi-analytic SN prescription, but the resulting remnant mass prescription differs. The model focuses more on the removal of the hydrogen-rich common envelope and the evolution steps before core collapse. They also explore the amount of fallback. Their models overestimate the BH masses by 10% if the BHs are not formed by direct collapse. They do not account for any convective boundary mixing beyond the core helium burning.
Schneider et al. (2023) uses the models in Schneider et al. (2021) and the stellar evolution code MESA (Paxton et al., 2011) to explore the black hole mass distribution and chirp mass distribution of BBH mergers. The ability of stars to explode in the delayed neutrino-driven SN mechanism is characterised by the compactness of the progenitor cores (O’Connor & Ott, 2011; Ertl et al., 2016). This depends on the mass and radius, and a high compactness indicates unsuccessful explosions and collapse into a BH. Thus, the chirp mass distribution is bimodal and the model predicts a lack of chirp masses in the range $9-13M_{\odot}$. Mass accretion can influence the final outcome of core collapse and final remnant mass. Schneider et al. (2024) studies models with MESA that allow for multiple mergers. They find BH masses in the pair instability SN gap, which can explain events such as GW190521. They note that for a more complete picture of stellar mergers, further detailed multi-dimensional, hydrodynamic simulations and a systematic exploration of the relevant parameter space is needed.
@@ 2.4 Summary
L2 point angular momentum loss linear fraction ($f_{\gamma}$) [0,0.25,0.5,0.75,1] 0.5
@@
Table 1: COMPAS hyper-parameters explored in this paper. Note that some prescriptions are only used in conjunction with others. The mass transfer fraction is only applied if the mass transfer accretion efficiency prescription is set to FIXED_FRACTION. The default is THERMAL. Similarly, the angular momentum loss linear fraction is only applied when the mass transfer angular momentum loss prescription is MACLEOD_LINEAR. The default is ISOTROPIC ($f_{\gamma}=0$). Refer to Tables 4,5 and 6 in Appendix C for all relevant prescriptions used for each model.
@@
    The binding energy formalism utilises a fixed $\alpha_{\rm CE}$, but is this a valid approach? Hirai & Mandel (2022) proposes a two-step CE process, which depends on the evolutionary stage of the donor and the companion’s mass. They find that the predicted post-CE separation is wider compared to the $\alpha_{\rm CE}$ formalism. This may pose a problem for BBHs merging within Hubble time. Picker et al. (2024) has recently implemented fits for the convective envelope mass and associated values of $\lambda$ to be used in conjunction with the two-step CE formalism. They find the effective combined value of $\alpha_{\rm CE}\lambda$ can range from $\sim 10^{-3}$ to $10^{3}$ when converted to the $\alpha$-formalism.
@@
    At masses $>50M_{\odot}$, Hurley et al. (2000) smoothly extrapolate the Pols et al. (1998) model grid up to $150M_{\odot}$, which is not well tested outside of the prescribed mass range. Significant improvement on this extrapolation method is far from trivial, as the evolution of high mass binaries is still very uncertain. This is reflected by the discrepancies between population synthesis codes in their stellar tracks (Agrawal et al., 2020, 2022a, 2022b).
@@
    There are many uncertainties in the mass transfer efficiency and angular momentum loss, which influence the Roche lobe response to mass loss. Uncertainties in how a donor star responds to mass loss, combined with the Roche lobe response uncertainties contributes to the lack of understanding around the mass transfer stability boundary as a function of donor mass and radius (Ge et al., 2010; Ge et al., 2015, 2020; Temmink et al., 2023).
@@
    Is the angular momentum loss occurring closer to the accretor or the L2 point? And should the implementation of $f_{\gamma}$ vary for degenerate and non-degenerate accretors? When performing the analysis in this paper, split functionality for AM loss from degenerate and non-degenerate accretors was yet to be implemented in COMPAS. Physically, it is more accurate for $f_{\gamma}$ to be implemented differently for the primary and secondary star, particular when the accretor is a compact object. It is not clear how this would impact our current results, where we have only considered the angular momentum loss near the non-degenerate object. Gallegos-Garcia et al. (2023) derive an analytic model for mass loss from the binary that is informed by observations of X-ray binary accretion disks and hydrodynamical simulations, which considers angular momentum loss from a disk wind around a compact object accretor. They find their model predicts more angular momentum loss compared to the standard case of isotropic re-emission and the orbital separation of a binary can be considerably affected by mass loss via a disk wind.
@@ 3 Predicting the GW merger rate for different COMPAS models
and $\psi(Z,t_{i})$ is the metallicity ($Z$) dependent star formation rate in the galaxy when the progenitor stars were born. $M_{\rm tot}^{\alpha}$ is a normalisation constant for a given model $\alpha$ (see Eq 10 in Rauf et al. 2023), which normalises the star formation rate based on the chosen stellar initial mass function, in such a way that it cancels out the dependence of the number of GW events on the number of binaries that we simulate $N_{\mathrm{binary}}$, and any lower or upper mass cuts we apply when running COMPAS. $\vartheta(t,t_{j},\Delta t_{j})$ is a counter for the number of BBH mergers in Shark snapshot, $j$. $\tau_{n}^{\alpha}(Z)$ is the coalescence time of each binary outputted from the model $\alpha$. If the merger time, $t=t_{i}+\tau_{n}^{\alpha}(Z)$, falls in the $j^{\mathrm{th}}$ bin, a count is added to that bin, otherwise it does not contribute. Overall, this equation encapsulates that the rate consists of two contributions, one from the Shark host galaxy and the other from population synthesis (in our case from COMPAS).
@@ 3.1 The effective factor for different models
In the second line we have substituted in Eq. 12 and use $\overline{\psi}$, $\overline{\theta}$ to acknowledge that this is the average of the star-formation history and Heaviside function over all galaxies that fall into the the bin $dZdt_{i}dt_{j}$ (i.e., we assume that all galaxies in this bin produce the same number of gravitational waves, which is strictly true only for infinitesimally small bins). We now can introduce a new population synthesis model, $\beta$, where the merger rate is unknown. If we have the merger rate from model $\alpha$, we can take a ratio of the two models and rearrange,
$\frac{d^{3}N_{\rm GW,k}^{\beta}}{dZdt_{i}dt_{j}}$ $=\frac{M_{\rm tot}^{\alpha}}{M_{\rm tot}^{\beta}}\frac{d^{3}N_{\rm GW}^{\alpha}}{dZdt_{i}dt_{j}}\frac{\overline{\sum_{n}\theta(t_{i}+\tau_{n}^{\beta}(Z,t_{i}),t_{j},\Delta t_{j})}}{\overline{\sum_{n}\theta(t_{i}+\tau_{n}^{\alpha}(Z,t_{i}),t_{j},\Delta t_{j})}}$
@@
We hence find that retaining the information from Shark is not required as the star formation history, $\overline{\psi}(Z,t_{i})$, and the galaxy distribution are independent of the population synthesis model. In practice we also expect that for most use cases $\frac{M_{\rm tot}^{\alpha}}{M_{\rm tot}^{\beta}}\approx 1$ since changes to the initial mass function between different literature values primarily affect stars that are too small to become binary black holes or neutron stars, although we retain it here for completeness. [^1] Note this factor also needs to be included if one were to rescale between models where different lower or upper mass limits have been used to draw stars from the Initial Mass Function (IMF), where a different mass ratio distribution for the initial binaries is assumed, or the overall binary fraction is changed The only caveat then is on the free choice of binning used to evaluate what we call the ‘effective factor’ $f^{\alpha\beta}_{\mathrm{eff}}$ between the two models — although the dependence on galaxy simulation has been absorbed into our fiducial $\frac{d^{3}N_{\rm GW}^{\alpha}}{dZdt_{i}dt_{j}}$, there is still the implicit assumption that the rate in each galaxy in the bin is the same. We find that this approximation remains valid over a range of metallicity binning choices, even for the quite wide bins we adopt as our default (100 un-evenly spaced bins for $\log Z\in[-7.0,-0.3]$). We show an example distribution of $f_{\rm eff}$ in these metallicity and merger time bins in Fig. 1 for a single choice of binary formation redshift of 2.5 (corresponding to $t_{i}\approx 2.62\rm Gyr$).
@@ 3.2 Generating the merger rate
    The starting point of our algorithm is a high-fidelity set of GW events simulated from a combination of Shark and COMPAS model $\alpha$. The detailed steps used for generating these are covered in Rauf et al. (2023) and so will not be repeated here. Briefly, this is done by cross-matching the galaxy metallicities to the binary population metallicities outputted from COMPAS. We inject this population into the Shark galaxies and track their evolution until they merge using Eq 16.
@@
    Having established our metallicity, birth time and merger time bins, we take our full list COMPAS binaries, read in the metallicity and evolution time, which is added to the birth time ($t_{\rm birth}+t_{\rm BBH}+t_{\rm coalescence}$), and add these to the corresponding bins. We apply this to both the fiducial and new model, take the ratio and multiply by the merger rate in (ii) to estimate the new merger rate, shown in Eq 18.
Figure 2: Volumetric BBH merger rate with remnant mass prescription from Mandel & Müller (2020) as a function of redshift. The blue line is the true merger rate from running COMPAS and Shark. The red dashed line is the merger rate rescaled with the effective factor calculated using the full COMPAS run containing $170\times 10^{6}$ binaries. The grey lines are the rescaled merger rate where the effective factor has been calculated using a unique sub-sample of $10^{6}$ binaries from the full COMPAS run. The green line is the average over 170 of these subsamples.
@@
$R_{\rm GW}^{\beta}(t_{\rm j})=\frac{1}{V_{\mathrm{sim}}}\sum_{i=0}^{170}\sum_{k=0}^{100}\frac{d^{3}N_{\rm GW}^{\beta}}{dZ_{k}dt_{i}dt_{j}},$ (19)
@@
For efficiency, we choose a sample $\tilde{N}_{\beta}^{\rm GW}$ of $10^{6}$ binaries from our $\beta$ COMPAS run to estimate the effective factor, while our fiducial model $\alpha$ has $170\times 10^{6}$. [^2] Note that we draw only $\sim 10^{6}$ unique masses for this process from the IMF, but each one is randomly assigned a different metallicity from Shark, resulting in $\sim 170\times 10^{6}$ unique combinations of initial masses and metallicities, and so $\sim 170\times 10^{6}$ unique binaries. The merger rate is reduced by a factor of 170 but this is compensated for by the normalisation $\frac{M_{\rm tot}^{\alpha}}{M_{\rm tot}^{\beta}}\sim 170$. Thus, this formalism removes the dependence on the sample size. We test the validity of our rescaling procedure and using only $10^{6}$ binaries for model $\beta$ by splitting one of our fiducial COMPAS runs from Rauf et al. (2023) (using the remnant mass model of Mandel & Müller 2020) into 170 by $10^{6}$ sub-samples. We then use these to estimate the rescaled merger rate for this model given $\alpha=\,$Fryer et al. (2012), and look at the spread in the merger rate. We show in Fig. 2 that overall the rescaling method using equal numbers of binaries for models $\alpha$ and $\beta$ recovers the true merger rate from our high-fidelity cross-matching, but also that using only $10^{6}$ binaries for model $\beta$ returns, on average, an unbiased estimate of the merger rate. The use of a smaller number of binaries in the rescaling procedure does introduce additional Poisson scatter in the merger rate, which we propagate through into our model comparison/evaluation (see Section 4.
@@
Figure 3: Volumetric rate plot as a function of redshift from Rauf et al. (2023) recreated with the effective factor scaling method. The error on the plots are the range of merger rates from sampling from each COMPAS run 170 times.

### v2

@@ 2.2 Chemically homogeneous evolution
Mandel & de Mink (2016) find that the CHE channel significantly contributes to the BBH merger rate ($10.5\pm 0.5$ $\rm Gpc^{-3}yr^{-1}$ at redshift 0). These systems typically have a total mass ranging between $50-110$ $M_{\odot}$. Marchant et al. (2016) explores the parameter space for forming CHE binaries. They find that at low metallicity, BBHs will merge within the age of the Universe due to weaker winds ($t_{\rm delay,min}=0.4$ Gyr for $Z_{\odot}/50$). There is strong evidence for increasing masses and mergers with decreasing metallicity (Belczynski et al., 2010; Bulik & Belczynski, 2010; Kowalska-Leszczynska et al., 2015; Spera et al., 2015; Neijssel et al., 2019), and for masses in the pair instability SN gap (van Son et al., 2020; Winch et al., 2024). [^2] Due to the production of electron positron pairs in the core of massive stars, the reduction of pressure followed by gravitational collapse leads to oxygen burning and results in a partial or complete explosion, which leaves behind no compact remnant (Fowler & Hoyle, 1964; Barkat et al., 1967). If these mergers were detected, they could probe the evolution of massive stars in the early Universe. In addition to calculating the merger rate and mass distribution for the CHE channel, du Buisson et al. (2020) finds that the cosmological merger rate ($5.8$ $\rm Gpc^{-3}yr^{-1}$ at redshift 0) can vary up to 40% for extreme deviations in the metallicity-specific SFR.
In COMPAS, there are two CHE modes, ‘optimistic’ and ‘pessimistic’, which refer to how a CH star evolves on the MS (Riley et al., 2021). In the ‘optimistic’ mode, if the star begins the main sequence with fast enough rotation to evolve as a CHE star, it is assumed the star continues rotating fast enough to remain CH throughout the MS and the spin is not checked. For the CHE mode ‘pessimistic’ the spin is checked at every timestep, and if it has slowed below the CH metallicity-dependent rotational frequency threshold, it is immediately converted to a regular MS star and evolve to the end of the MS (Team COMPAS: Riley, J. et al., 2022). We choose to study both variations in COMPAS, as well as no CHE, to determine if the CHE can be ruled out as a formation channel for GW sources.
@@ 2.3 Remnant mass prescription
GWs can be used to infer the underlying mass distribution of BBHs, which depends on the metallicity and the type of explosion mechanism. Their formation can be described in three stages; stellar collapse and bounce, convective engine, and post-explosion fallback. Massive stars undergo a series of successful burning stages, building an iron core until electron degeneracy pressure can no longer support this. Electron capture and dissociation of the core elements into alpha particles accelerates the core collapse. This collapse is halted when the core reaches nuclear densities and nuclear forces cause the core to bounce and emit a shock outwards. This results in a mixing of the hot and cold layers in the star, known as convection. The convective engine provides the means to convert potential energy released in the collapse into explosion energy in the form of a supernova (Fryer et al., 2012; Fryer et al., 2021). The commonly accepted supernova explosion mechanism is neutrino driven, where some of the neutrinos emitted during the core collapse are reabsorbed, powering an explosion (Janka et al., 2007; Marek & Janka, 2009; Janka, 2012, 2017). We choose to explore all remnant mass prescriptions available in the COMPAS version used in this paper, and describe them below.
Fryer et al. (2012) provide analytic prescriptions for estimating the remnant mass in population synthesis codes, based on at the time recent understanding of supernova and gamma-ray burst explosions. In their model/fit, the remnant mass depends on the core, fallback mass and explosion mechanism. They explore two neutrino-driven explosion mechanisms– the rapid convection, which leads to explosion in the 250 ms after bounce, and the delayed convection, which can occur over a longer timescale. Fryer et al. (2022) altered the remnant mass prescription from Fryer et al. (2012) to study the growth rate of convection. They introduce a new parameter to describe the mixing growth time for convective instabilities in their remnant mass models, denoted $f_{\rm mix}$. A fast growth time, $f_{\rm mix}=4.0$, and a slow growth time, $f_{\rm mix}=0.5$, correspond to the rapid and delayed models respectively from Fryer et al. (2012). The growth time determined the supernova explosive energy (see Fig. 4 in Fryer, 2015). Whether a star undergoes a successful or failed supernova provides clues on the final remnant object. However, there are uncertainties surrounding these mechanisms (Mapelli, 2021). GW detection from a core collapse SN can be used to distinguish between explosion mechanisms (Powell et al., 2016, 2023)
@@
Schneider et al. (2021) also applies the Mandel & Müller (2020) semi-analytic SN prescription, but the resulting remnant mass prescription differs. The model focuses more on the removal of the hydrogen-rich envelope of the progenitor star and the evolution steps before core collapse. They also explore the amount of fallback. Their models overestimate the BH masses by 10% if the BHs are not formed by direct collapse. They do not account for any convective boundary mixing beyond the core helium burning.
Schneider et al. (2023) uses the models in Schneider et al. (2021) and the stellar evolution code MESA (Paxton et al., 2011) to explore the black hole mass distribution and chirp mass distribution of BBH mergers. The ability of stars to explode in the delayed neutrino-driven SN mechanism is characterised by the compactness of the progenitor cores (O’Connor & Ott, 2011; Ertl et al., 2016). This depends on the mass, radius and a nuclear reaction network, and a high compactness indicates unsuccessful explosions and collapse into a BH. Thus, the chirp mass distribution is bimodal and the model predicts a lack of chirp masses in the range $9-13M_{\odot}$. Mass accretion can influence the final outcome of core collapse and final remnant mass. Schneider et al. (2024) studies models with MESA that allow for multiple mergers. They find BH masses in the pair instability SN gap, which can explain events such as GW190521. They note that for a more complete picture of stellar mergers, further detailed multi-dimensional, hydrodynamic simulations and a systematic exploration of the relevant parameter space is needed.
@@ 2.4 Summary
L2 point AM loss linear fraction ($f_{\gamma}$) [0,0.25,0.5,0.75,1] 0.5
@@
Table 1: COMPAS hyper-parameters explored in this paper. Note that some prescriptions are only used in conjunction with others. The mass transfer fraction is only applied if the mass transfer accretion efficiency prescription is set to FIXED_FRACTION. The default is THERMAL. Similarly, the AM loss linear fraction is only applied when the mass transfer AM loss prescription is MACLEOD_LINEAR. The default is ISOTROPIC ($f_{\gamma}=0$). Refer to Tables 4,5 and 6 in Appendix C for all relevant prescriptions used for each model.
@@
    The binding energy formalism utilises a fixed $\alpha_{\rm CE}$, but is this a valid approach? The parametrisation of $\alpha_{\rm CE}$ is a point of contention, particularly when additional energy sources are available. It is also not clear how these would scale with the orbital separation and mass, so the $\alpha$-formalism is not an accurate representation of CE ejection. Hirai & Mandel (2022) proposes a two-step CE process, which depends on the evolutionary stage of the donor and the companion’s mass. They find that the predicted post-CE separation is wider compared to the $\alpha_{\rm CE}$ formalism. This may pose a problem for BBHs merging within Hubble time. Picker et al. (2024) has recently implemented fits for the convective envelope mass and associated values of $\lambda$ to be used in conjunction with the two-step CE formalism. They find the effective combined value of $\alpha_{\rm CE}\lambda$ can range from $\sim 10^{-3}$ to $10^{3}$ when converted to the $\alpha$-formalism.
@@
    At ZAMS masses $>50M_{\odot}$, Hurley et al. (2000) smoothly extrapolate the Pols et al. (1998) model grid up to $150M_{\odot}$, which is not well tested outside of the prescribed mass range. We find the contribution of $M_{\rm ZAMS}>50M_{\odot}$ to BBH mergers in COMPAS is non-negligible. This mass range could be significant as LVK detects more GW events in the pair instability mass gap (Winch et al., 2024). Significant improvement on this extrapolation method is far from trivial, as the evolution of high mass binaries is still very uncertain. This is reflected by the discrepancies between population synthesis codes in their stellar tracks (Agrawal et al., 2020, 2022a, 2022b).
@@
    There are many uncertainties in the mass transfer efficiency and AM loss, which influence the Roche lobe response to mass loss. Uncertainties in how a donor star responds to mass loss, combined with the Roche lobe response uncertainties contributes to the lack of understanding around the mass transfer stability boundary as a function of donor mass and radius (Ge et al., 2010; Ge et al., 2015, 2020; Temmink et al., 2023).
@@
    Is the AM loss occurring closer to the accretor or the L2 point? And should the implementation of $f_{\gamma}$ vary for degenerate and non-degenerate accretors? When performing the analysis in this paper, split functionality for AM loss from degenerate and non-degenerate accretors was yet to be implemented in COMPAS. Physically, it is more accurate for $f_{\gamma}$ to be implemented differently for the primary and secondary star, particular when the accretor is a compact object. It is not clear how this would impact our current results, where we have only considered the AM loss near the non-degenerate object. Gallegos-Garcia et al. (2023) derive an analytic model for mass loss from the binary that is informed by observations of X-ray binary accretion disks and hydrodynamical simulations, which considers AM loss from a disk wind around a compact object accretor. They find their model predicts more AM loss compared to the standard case of isotropic re-emission and the orbital separation of a binary can be considerably affected by mass loss via a disk wind.
@@
    The COMPAS version used in this paper does not consider tidal evolution, but it is believed to play a significant role in the evolution of massive stars (Team COMPAS: Riley, J. et al., 2022). Tidal forces can bring the compact object into the envelope of the secondary which is now a giant star. This implies that COMPAS underestimates the number of BH-giant system mergers, while overestimating the number of BBH mergers (Grichener, 2023). When performing the analysis in this paper, the newest version of COMPAS implemented tides (Kapil et al., in prep.).
  8. 8.
@@ 3 Predicting the GW merger rate for different COMPAS models
[…]
@@ 3.1 The effective factor for different models
[…]
$\frac{d^{3}N_{\rm GW}^{\beta}}{dZdt_{i}dt_{j}}$ $=\frac{M_{\rm tot}^{\alpha}}{M_{\rm tot}^{\beta}}\frac{d^{3}N_{\rm GW}^{\alpha}}{dZdt_{i}dt_{j}}\frac{\overline{\sum_{n}\theta(t_{i}+\tau_{n}^{\beta}(Z,t_{i}),t_{j},\Delta t_{j})}}{\overline{\sum_{n}\theta(t_{i}+\tau_{n}^{\alpha}(Z,t_{i}),t_{j},\Delta t_{j})}}$
@@
We hence find that retaining the information from Shark is not required as the star formation history, $\overline{\psi}(Z,t_{i})$, and the galaxy distribution are independent of the population synthesis model. In practice we also expect that for most use cases $\frac{M_{\rm tot}^{\alpha}}{M_{\rm tot}^{\beta}}\approx 1$ since changes to the IMF between different literature values primarily affect stars that are too small to become binary black holes or neutron stars, although we retain it here for completeness. [^5] Note this factor also needs to be included if one were to rescale between models where different lower or upper mass limits have been used to draw stars from the IMF, where a different mass ratio distribution for the initial binaries is assumed, or the overall binary fraction is changed The only caveat then is on the free choice of binning used to evaluate what we call the ‘effective factor’ $f^{\alpha\beta}_{\mathrm{eff}}$ between the two models — although the dependence on galaxy simulation has been absorbed into our fiducial $\frac{d^{3}N_{\rm GW}^{\alpha}}{dZdt_{i}dt_{j}}$, there is still the implicit assumption that the rate in each galaxy in the bin is the same. We find that this approximation remains valid over a range of metallicity binning choices, even for the quite wide bins we adopt as our default (100 un-evenly spaced bins for $\log Z\in[-7.0,-0.3]$). We show an example distribution of $f_{\rm eff}$ in these metallicity and merger time bins in Fig. 1 for a single choice of binary formation redshift of 2.5 (corresponding to $t_{i}\approx 2.62\rm Gyr$).
@@ 3.2 Generating the merger rate
    The starting point of our algorithm is a high-fidelity set of GW events simulated from a combination of Shark and COMPAS model $\alpha$. The detailed steps used for generating these are covered in Rauf et al. (2023) and so will not be repeated here. Briefly, this is done by cross-matching the galaxy metallicities to the binary population metallicities outputted from COMPAS. We inject this population into the Shark galaxies and track their evolution until they merge using Eq. 16.
@@
    Having established our metallicity, birth time and merger time bins, we take our full list of COMPAS binaries, read in the metallicity and evolution time, which is added to the birth time ($t_{\rm birth}+t_{\rm BBH}+t_{\rm coalescence}$), and add these to the corresponding bins. We apply this to both the fiducial and new model, take the ratio and multiply by the merger rate in (ii) to estimate the new merger rate, shown in Eq. 18.
Figure 2: Volumetric BBH merger rate with remnant mass prescription from Mandel & Müller (2020) as a function of redshift. The blue line is the Rauf et al. (2023) merger rate from running COMPAS and Shark. The red dashed line is the merger rate rescaled with the effective factor calculated using the full COMPAS run containing $170\times 10^{6}$ binaries. The grey lines (whose overlapping nature visually appears as a shaded black region) are the rescaled merger rate where the effective factor has been calculated using 170 unique sub-samples of $10^{6}$ binaries from the full COMPAS run. The green line is the average over these sub-samples.
@@
$R_{\rm GW}^{\beta}(t_{j})=\frac{1}{V_{\mathrm{sim}}}\sum_{i=0}^{170}\sum_{k=0}^{100}\frac{d^{3}N_{\rm GW}^{\beta}}{dZ_{k}dt_{i}dt_{j}},$ (19)
@@
For efficiency, we choose a sample $\tilde{N}_{\rm GW}^{\beta}=10^{6}$ binaries from our $\beta$ COMPAS run to estimate the effective factor while using the full sample for our fiducial model, $\alpha$, where $N_{\rm GW}^{\alpha}=170\times 10^{6}$. [^6] Note that we draw only $\sim 10^{6}$ unique masses for this process from the IMF, but each one is randomly assigned a different metallicity from Shark, resulting in $\sim 170\times 10^{6}$ unique combinations of initial masses and metallicities, and so $\sim 170\times 10^{6}$ unique binaries. The merger rate is reduced by a factor of 170 but this is compensated for by the new normalisation in Eq. 18, $\frac{M_{\rm tot}^{\alpha}}{M_{\rm tot}^{\beta}}\sim 170$. Thus, this formalism removes the dependence on the sample size. We test the validity of our rescaling procedure and using only $10^{6}$ binaries for model $\beta$ by splitting one of our fiducial COMPAS runs from Rauf et al. (2023) (using the remnant mass model of Mandel & Müller 2020) into 170 by $10^{6}$ sub-samples. We then use these to estimate the rescaled merger rate for this model given $\alpha=\,$Fryer et al. (2012), and look at the spread in the merger rate. We show in Fig. 2 that overall the rescaling method using equal numbers of binaries for models $\alpha$ and $\beta$ recovers the Rauf et al. (2023) merger rate from our high-fidelity cross-matching, but also that using only $10^{6}$ binaries for model $\beta$ returns, on average, an unbiased estimate of the merger rate. The use of a smaller number of binaries in the rescaling procedure does introduce additional Poisson scatter in the merger rate, which we propagate through into our model comparison/evaluation (see Section 4).
@@
Figure 3: Volumetric rate plot as a function of redshift from Rauf et al. (2023) recreated with the effective factor scaling method. The error on the plots are the range of merger rates from sampling from each COMPAS run 170 times. We use a logarithmic scale for the y-axis, as the merger rates for the models spans several orders of magnitude.